Please find:
the distance (-2,3,1) and (4,5,-1)
and
midpoint (-2,3,1) and (4,5,-1)
Thanks

Answer by venugopalramana(3286)

Please find:
the distance(X1,Y1,Z1)= (-2,3,1) and (X2,Y2,Z2)=(4,5,-1)
DISTANCE = SQRT[(X2-X1)^2+(Y2-Y1)^2+(Z2-Z1)^2]
=SQRT[6^2+2^2+2^2]=SQRT(44)=2SQRT(11)
and
midpoint (X1,Y1,Z1)=(-2,3,1) and (X2,Y2,Z2)=(4,5,-1)
MID POINT = [(X1+X2)/2,(Y1+Y2)/2,(Z1+Z2)/2]
= [2/2,8/2,0/2]=(1,4,0)
Thanks